OpenStudy (anonymous):

The planets move around the sun in elliptical orbits with the sun at one focus. The point in the orbit at which the planet is closest to the sun is called perihelion, and the point at which it is farthest is called aphelion. These points are the vertices of the orbit. A planet's distance from the sun is 224,000,000 km at perihelion and 232,000,000 km at aphelion. Find an equation for the planet's orbit. (Place the origin at the center of the orbit with the sun on the x-axis.

OpenStudy (anonymous):

Taking astronomy?

OpenStudy (anonymous):

lol no this is supposed to be "college algebra"

OpenStudy (anonymous):

well, you can take the average of the two distances and make that the center point I suppose. then the sun would be off to the left on the x axis and you'd have an ellipse and know where it crosses the x-axis, but I can't see how to find the y intercepts of the elipse.. like I said, -need to study these much more..

OpenStudy (anonymous):

ok so then the center point is 228,000,000. Wouldn't the center point have something to do with the origin?

OpenStudy (anonymous):

that's what I'm trying.. yea, the center is at 228,000,000 and so the sun is off to the left by 4,000,000 right?? :/

OpenStudy (anonymous):

yeah that would seem right

OpenStudy (anonymous):

I'm confused.. I'm not sure I'm doing it right...

OpenStudy (anonymous):

yea that's right.

OpenStudy (anonymous):

yeah it's logical

OpenStudy (anonymous):

so, umm what was the equation for an ellipse? uummmm.. \{x^2}{a^2} + {y^2}{b^2} = 1.

OpenStudy (anonymous):

division there between them..

OpenStudy (anonymous):

You can try reposting this to see if someone else bites too. I'm just trying..

OpenStudy (anonymous):

yeah i'm pretty sure that's the equation for an ellipse. I'll try that

OpenStudy (phi):

You can look at http://en.wikipedia.org/wiki/Ellipse

OpenStudy (anonymous):

is "a" the distance from 0 to the x intercept? and "b" the distance to the y intercept?

OpenStudy (anonymous):

I think I know how to do it... by using this focus formula.... we can find b... Focus The distance from the center C to either focus is f = ae, which can be expressed in terms of the major and minor radii: f = sqrt{a^2-b^2}.

OpenStudy (anonymous):

4,000,000 = root(a^2 - b^2)

OpenStudy (anonymous):

solve for b

OpenStudy (anonymous):

because we know a...

OpenStudy (anonymous):

what was a again?

OpenStudy (anonymous):

319,000,000 "the difference of the aphilion and the distance from the sun to the origin... 232,000,000-4,000,000

OpenStudy (anonymous):

wait I'm on crack..

OpenStudy (anonymous):

a = 228000000.. lol

OpenStudy (anonymous):

(x^2)/(2.28x10^8)^2 + (y^2)/b^2 = 1 were b is to be solved in this... 4,000,000 = root(a^2 - b^2) That's the best I can do right now and am not totally sure it's right...

OpenStudy (anonymous):

sorry i didn't respond for a while... i had to reboot my computer lol. thanks for helping

OpenStudy (anonymous):

NP i'm sorta going back and forth between this and other things anyway..

OpenStudy (anonymous):

but I think that's the one...so after solving for b I get... b = 227964909.6 , so sub that in and if we haven't made any mistakes it should be ...close.. lol I would run through it again though to be sure if you care to..

OpenStudy (anonymous):

i just need the equation for the planet's orbit

OpenStudy (anonymous):

(x^2)/(2.28x10^8)^2 + (y^2)/(2.27^8)^2 = 1 That's what I got...

OpenStudy (anonymous):

ok thanks